13. If the events X and Y are independent and P(X) = .4 and P(Y) = 5, what does P(XY) equal? (2) What is the probability of X or Y. (2) What is the conditional probability of Y given X? (2) If the events X and Y are independent and P(X) = .4 and P(Y)-3, what is the conditional probability of Y given X?(2) #4
Answer #1

3.

P(X) = .4

P(Y) = .5

P(XY) = P(X) * P(Y) = .4*.5 = .20

P(X or Y) = P(X) + P(Y) = .4 + .5 = .9

P(Y/X) = P(XY)/P(X) = .2/.4 = .5

4.

P(Y/X) = P(XY)/P(X) = (.4*.3)/.4 = .3